OpenStudy (yourboydusty):

What is the slope of a line that crosess through the points (-4,-5) and (6,8)?

OpenStudy (mecharv):

Do you know the two point formula?

OpenStudy (yourboydusty):

yes but i got it wrong somehow??? Slope = rise/run = y2-y1/x2-x1

OpenStudy (mecharv):

Okay, that is correct. Just substitute the points in the formula and thats your answer. Slope =y2−y1/x2−x1 =8−(−5)/(6)−(−4)=8+5/6+4=13/10

OpenStudy (yourboydusty):

thats alot of numbers.....

OpenStudy (mecharv):

Oh not at all, just give me a second

OpenStudy (yourboydusty):

thank you srry i get confused very easy :(

OpenStudy (mecharv):

First see what are the two points given to you. In this case \[\left( -4,-5 \right)\] and \[\left( 6,8 \right)\] Now remember for this case like in alphabets after x comes y, so in an ordered pair, \[\left( x_{1},y_{1} \right)\] and \[\left( x_{2},y_{2} \right)\] Now just identify \[x1,y1,x2,y2 \] by directly comparing the two. In this case \[x_{1}=-4, y_{1}=-5, x_{2}=6, y_{2}=8\]

OpenStudy (yourboydusty):

okay thank you....could u give me the answer so i can make sure i got it right

OpenStudy (mecharv):

Now use your formula, i.e, \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] \[m=\frac{ 8-(-5) }{ 6-(-4) }= \frac{ 8+5 }{ 6+4 }=\frac{ 13 }{ 10 }\]

OpenStudy (yourboydusty):

yay i did get it right!!!\

OpenStudy (mecharv):

You understood right?

OpenStudy (yourboydusty):

yes thyank you

OpenStudy (callielovesyhuu):

-.-